Change the following numbers into there binary equivalents..... 19,39 and 77
Answer by stanbon(75887)

Change the following numbers into there binary equivalents..... 19,39 and 77
-----
19 = 16+2+1 = 2^4 + 2^1 + 1 = 10011
-----
39 = 32+4+2+1 = 2^5+2^2+2^1+1 = 100111
------
77 = 64+8+4+1 = 2^6+2^3+2^2+1 = 1001101
---------------
Cheers
Stan H.
----------